# Self-Taught Agentic Long-Context Understanding

### v1

@@ 6.1 Tasks and Metrics
We use GPT-4o as the judge for answer correctness, with the prompt template shown in appendix E. We report accuracies for all datasets.
@@ Appendix B Short Context Performance
Appendix C Detailed Results on Seven Benchmark Tasks
As shown in table 9table 10table 11table 12table 13table 14 and table 15, we evaluate the long-context performance across seven tasks: HotpotQA, Natural Questions, TriviaQA, PopQA, NarrativeQA, InfiniteBench QA and InfiniteBench Multiple-Choice. AgenticLU provides significant improvement for all tasks, especially for those that require multi-hop reasoning such as HotPotQA.
@@ Appendix C Detailed Results on Seven Benchmark Tasks
Appendix D Chain-of-Clarifications Workflow
@@
Appendix E Evaluation Template

### v2

@@ 6.1 Tasks and Metrics
We use GPT-4o as the judge for answer correctness, with the prompt template shown in appendix F. We report accuracies for all datasets.
@@ 8 Conclusion
Acknowledgement
Our work is sponsored in part by NSF CAREER Award 2239440, NSF Proto-OKN Award 2333790, Sponsored Research Projects from companies like Cisco and eBay, as well as generous gifts from Google, Adobe, and Teradata. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and should not be interpreted as necessarily representing the views, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for government purposes not withstanding any copyright annotation hereon.
@@ Appendix B Short Context Performance
Appendix C Agentic Workflow without Training
We conducted an additional experiment for prompting only with our agentic workflow, and we find that with an option to generate clarifications the model does get better on multi-hop QA questions (HotPotQA), but it is generally difficult for the base model to point to the correct paragraphs directly without any training, hence often resulting in same or slightly worse performance.
Results are listed in table 9, the context length is 128K tokens.
Table 9: Performance comparison across different QA benchmarks with 128K token context length, with or without training.
Model HotpotQA NaturalQ PopQA TriviaQA Avg
Llama-3.1-8B 40.0 56.1 56.1 80.6 58.2
Llama-3.1-8B + Prompting Only 53.3 56.7 51.6 72.8 58.6
AgenticLU-8B 71.1 77.8 65.5 88.3 75.7
Appendix D Detailed Results on Seven Benchmark Tasks
As shown in table 10table 11table 12table 13table 14table 15 and table 16, we evaluate the long-context performance across seven tasks: HotpotQA, Natural Questions, TriviaQA, PopQA, NarrativeQA, InfiniteBench QA and InfiniteBench Multiple-Choice. AgenticLU provides significant improvement for all tasks, especially for those that require multi-hop reasoning such as HotPotQA.
@@ Appendix D Detailed Results on Seven Benchmark Tasks
Appendix E Chain-of-Clarifications Workflow
@@
Appendix F Evaluation Template